LR Engines stock is selling for $42.39 a share, has an ROE of 14.3 percent, and a dividend payout ratio of 35 percent. The next expected dividend is $1.62 a share. What is the cost of equity for this firm?

1. 12.86 percent

2. 13.12 percent

3. 13.47 percent

4. 12.52 percent
Answer #1

Growth =b*r =(1-0.35)*14.3% =9.295%

Cost of equity =D1/P0+g =1.62/42.39+9.295% =13.12%